Calculate osmotic pressure at 273k of a 5% solution of urea.(mole mass of urea=60)

Grade:12

1 Answers

Arun
6 years ago
Dear Student
 
=1/12Hence Concentration(C) of urea (in terms of Molarity)= (mole of urea(n)/Volume of solution)*1000={(1/12)/100}*1000=10/12Hence Osmotic pressure=1*(10/12)*0.082*272 atm=18.59 atm-1T=temperature (in Kelvin)i=Van’t-Hoff factor(=1 for non-electrolyte)5% urea solution means 5g urea is present in 100ml of solution.mole of urea=weight given/Molecular weight of urea= 5g/60gmol-1K-1Osmotic presuure=i*C*R*T,where, C=concentration of solute(in terms of Molarity)R= Gas constant=0.082 L(atm)(mol) RegardsArun (askIITians forum expert)